Austin Systems Inc. is expected to pay a $1.61 dividend at year end, the dividend is expected to grow at a constant rate of 2.7% a year, and the common stock currently sells for $23.75 a share. The before-tax cost of debt is 7.50%, and the tax rate is 40%. The target capital structure consists of 45% debt and 55% common equity. What is the company’s WACC if all the equity used is from retained earnings? Enter the answer as a decimal with four places of precision (i.e. 0.1234).
Answer #1

Cost of equity=(D1/Current price)+Growth rate

=(1.61/23.75)+0.027

=9.47894737%

Cost of debt after-tax=7.5*(1-tax rate)

=7.5*(1-0.4)=4.5%

WACC=Respective costs*Respective weight

=(9.47894737*0.55)+(0.45*4.5)

=0.0724(Approx).